OpenStudy (anonymous):

what is this as a fraction in simplest form??NEED HELP MAJORLYY! 2 2 --- - ____ 3 -7

OpenStudy (anonymous):

if u could please help me that would be awesome!!

OpenStudy (anonymous):

Take the LCM of the denominators i.e. -21 now rewrite the fractions 2/3 = -14/-21 and 2/-7 = 6/-21 now, 14 6 14-6 --- - ---- = -------- -21 -21 -21 = 8/-21

OpenStudy (anonymous):

it said that was incorrect

OpenStudy (anonymous):

it said it is 20/21

OpenStudy (anonymous):

Take the LCM of the denominators i.e. -21 now rewrite the fractions 2/3 = -14/-21 and 2/-7 = 6/-21 now, -14 6 -14-6 --- - ---- = -------- -21 -21 -21 = -20/-21 = 20/21

OpenStudy (anonymous):

okayy! what ab 9/-5 + 2/9

OpenStudy (anonymous):

same, Take the LCM of the denominators i.e. -45 now rewrite the fractions 81/-45 + -10/-45 71/-45

OpenStudy (anonymous):

it said the answer is -62/45

OpenStudy (anonymous):

then check the fractions in the question u can compare the answer just by pulging in the calc, u will get decimal answers

OpenStudy (anonymous):

wolframalph.com